A train is travelling23m/s its speed expressed in km is
1
Expert's answer
2020-08-11T08:27:33-0400

"60 sec = 1 min"

"60 min =1 hr"

"1 km = 1000 m"

"\frac{23m}{s}*\frac{60s}{1min}*\frac{1km}{1000m}*\frac{60min}{1hr}
=82.8 km/hr"